# Biomedical Open Source Software: Crucial Packages and Hidden Heroes

### v2

@@ 3 Results
Figure 3: Distribution of packages by Katz centrality and counts of their mentions in papers. Katz centrality is calculated for an unweighted graph, for a weighted graph with all nodes, or just for the largest connected cluster (LCC). In the calculations, we assumed $\beta=1$.
@@ 4 Discussion
The absence of cycles in the network of open-source software packages used in our sample of biomedical science is quite interesting. It suggests a more robust design structure in the universe of scientific software than in the general software world. The lack of cycles also improves our analysis with Katz centrality, as it excludes feedback mechanisms that can artificially inflate the centrality of packages on a loop. Packages with low mentions and high Katz centrality are therefore even more critical, as they represent essential dependencies of packages that enable large volumes of research without receiving direct attention. These are exactly the “Nebraskan” packages referenced by Munroe (2020) that we set out to find. Figure 3 seems to show one clear example of such a package: velvet, a PyPI package with a mention count of $197$, but a Katz centrality value calculated as $0.105$ for the largest connected cluster and as $0.053$ for the complete graph. Interestingly, for the unweighted representation, this package has a low centrality of $0.007$.
@@
The difference between the unweighted and weighted version of our analysis, in Figure 3 can help us parse different subtypes of these outliers. For example, the velvet package discussed above is central solely because it is mentioned by one highly cited paper (and dozens of less cited ones). This strategy is not central in the unweighted version of the analysis, since citation counts are ignored. Central packages under this representation therefore also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
Altogether, our paper tries to quantify impact as importance, and to put some numbers behind the intuition. Of course, this is by itself a perilous process: there could be various flavors of both impact and importance, and people may reasonably disagree about the details. Our contribution is to propose a certain quantification, which agrees with our intuitive understanding.
Besides the possible problems with the exact definition of importance, our work has a number of limitations:
@@
    The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions Dataset (see the discussion in (Istrate et al., 2022a)). Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms (Druskat et al., 2024). Notably, this affects the above-mentioned “Nebraskan” package velvet, which is linked to a Python package of the same name that provides signal processing and communications algorithms in Python (available on PyPI [^1] https://pypi.org/project/velvet/). Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short read assembly using de Bruijn graphs in genomics, as described in (Zerbino and Birney, 2008) and implemented mostly in C (Zerbino et al., 2014). Sometimes an open source package shares a name with a proprietary software (e. g. PRISM, PACE), which leads to confusion and incorrect usage counts.
@@
    Our approach is limited to the open source packages included into one of the chosen package managers. This excludes proprietary software like Excel or GraphPad Prism, and open source non-packaged software like Gephi (Bastian et al., 2009). The number of missing packages may be estimated from Table 14 of (Istrate et al., 2022a) which gives the coverage by GitHub at 64.39%, the coverage by CRAN at 8.36%, the coverage by PyPI at 5.86%, and the coverage by Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages – perhaps growing in number – are not hosted by GitHub, their fraction remains anecdotally low. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e. between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open source non-package software, it may be feasible to identify source repository URLs and then to identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services); nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers222as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies.
@@
    In our version of the dependency chart, we always used the latest version of any package. In reality, dependencies change between the versions. A study of development logs often reveals messages like “deleted the dependence on XYZ” or “added the dependence on ABC”. These changes often make the work of package management software very difficult. A software dependency graph is a living network, constantly evolving, with links added and deleted. In this study, we effectively collapse the time, making an (imperfect) snapshot of the long movie.
@@
    Our method captures the off the shelf software packages used for the papers. It does not capture ad hoc software written specifically for the given paper, which may be contained in the code accompanying the papers. This code may load software libraries, and thus add software dependencies absent in our graph.
@@
    The dependencies reflected in the package managers do not necessarily reflect the actual usage of the software. A library may be loaded, but not used for the actual execution path used in the given paper. A more reliable method would be to reproduce the analysis done in each paper and capture the actual library calls—which is probably prohibitively difficult for the number of papers covered.
@@
    Many Python and R packages ultimately depend on libraries outside their ecosystems (most often C and Fortran libraries). These dependencies are not easily captured by the package management software, and thus are outside of our analysis.
@@
    Our dependency graph does not include the concept of alternatives, when package $A$ depends on the functionality that can be provided either by package $B_{1}$, or package $B_{2}$. Some package managers like Debian’s apt have the possibility of specifying alternatives, but this is not a common feature of package managers.
@@
    Different communities place different emphasis on infrastructure elements such as tests. This may lead to the over-representation of testing infrastructure packages in some cases, for example in the Python ecosystem. We do not try to attempt to decide which elements of the infrastructure are “actually important”, instead relying on the judgment of the community of users and developers—while recognizing that such judgment may be different across the landscape.
@@
    There are several different Python package managers with greatly overlapping dependency graphs. We used PyPI for this study. Some alternatives like pip and Conda are also worth investigating.

### v3

@@ 3 Results
Fig 4: Distribution of packages by Katz centrality and counts of their mentions in papers. Katz centrality is calculated for an unweighted graph, for a weighted graph with all nodes, or just for the largest connected cluster (LCC) for each ecosystem. In the calculations, we assumed $\beta=1$.
Table 3: Software quadrants, unweighted graph
Software Ecosystem Mentions Centrality
Count Percentile Value Percentile
“Pasteur”
PRISMA cran 15 7971579715\,79715 797 $0.9997$ 0.277 620.277620.277\,620.277 62 $1$
DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.198 880.198880.198\,880.198 88 $0.9997$
ggplot2 cran 14 4411444114\,44114 441 $0.9995$ 0.234 070.234070.234\,070.234 07 $0.9999$
pymol pypi 14 6041460414\,60414 604 $0.9996$ 0.191 010.191010.191\,010.191 01 $0.9996$
PERMANOVA cran 21 9602196021\,96021 960 $1$ $0.1219$ $0.9992$
limma bioconductor 13 4511345113\,45113 451 $0.9993$ 0.217 930.217930.217\,930.217 93 $0.9998$
SAM cran 13 0481304813\,04813 048 $0.9992$ 0.144 250.144250.144\,250.144 25 $0.9994$
edgeR bioconductor 12 9231292312\,92312 923 $0.9991$ 0.171 370.171370.171\,370.171 37 $0.9995$
lme4 cran 10 5101051010\,51010 510 $0.999$ 0.131 910.131910.131\,910.131 91 $0.9993$
WGCNA cran 13 9151391513\,91513 915 $0.9994$ 0.097 210.097210.097\,210.097 21 $0.9988$
vegan cran 10 0121001210\,01210 012 $0.9989$ 0.121 210.121210.121\,210.121 21 $0.9991$
tophat pypi $7642$ $0.9987$ $0.1184$ $0.999$
STAR cran $9091$ $0.9988$ 0.110 520.110520.110\,520.110 52 $0.9989$
Popular
GSVA bioconductor $2937$ $0.9969$ 0.001 080.001080.001\,080.001 08 $0.361$
MAST bioconductor $2091$ $0.995$ 0.001 080.001080.001\,080.001 08 $0.361$
FSA cran $499$ $0.9782$ 0.001 080.001080.001\,080.001 08 $0.361$
Boruta cran $441$ $0.9754$ 0.001 080.001080.001\,080.001 08 $0.361$
MixSIAR cran $377$ $0.9714$ 0.001 170.001170.001\,170.001 17 $0.4989$
DRIMSeq bioconductor $335$ $0.9678$ 0.001 170.001170.001\,170.001 17 $0.4989$
netDx bioconductor $174$ $0.9419$ 0.001 070.001070.001\,070.001 07 00
COCOA bioconductor $162$ $0.9378$ 0.001 170.001170.001\,170.001 17 $0.4989$
emoji pypi $161$ $0.9374$ 0.001 070.001070.001\,070.001 07 00
aviso pypi $160$ $0.937$ 0.001 170.001170.001\,170.001 17 $0.4989$
SiZer cran $159$ $0.9366$ 0.001 170.001170.001\,170.001 17 $0.4989$
searchlight pypi $137$ $0.9273$ 0.001 070.001070.001\,070.001 07 00
“Nebraska”
vctrs cran $2$ $0.2516$ 0.024 380.024380.024\,380.024 38 $0.9928$
withr cran $2$ $0.2516$ 0.024 380.024380.024\,380.024 38 $0.9928$
isoband cran 00 00 0.024 380.024380.024\,380.024 38 $0.9928$
gss cran $7$ $0.4991$ 0.012 120.012120.012\,120.012 12 $0.9849$
SuppDists cran $3$ $0.3287$ 0.005 670.005670.005\,670.005 67 $0.9658$
lfda cran $1$ $0.1232$ $0.0049$ $0.9581$
ggtext cran $7$ $0.4991$ 0.004 820.004820.004\,820.004 82 $0.957$
texttable pypi $1$ $0.1232$ 0.004 780.004780.004\,780.004 78 $0.9569$
affyio bioconductor $5$ $0.4326$ 0.004 730.004730.004\,730.004 73 $0.956$
dnaio pypi $1$ $0.1232$ 0.004 720.004720.004\,720.004 72 $0.9558$
requests pypi $7$ $0.4991$ 0.004 190.004190.004\,190.004 19 $0.9499$
tifffile pypi $4$ $0.3863$ 0.004 090.004090.004\,090.004 09 $0.9488$
Table 4: Software quadrants, weighted graph (full)
Software Ecosystem Mentions Centrality
Count Percentile Value Percentile
“Pasteur”
PRISMA cran 15 7971579715\,79715 797 $0.9997$ 0.128 620.128620.128\,620.128 62 $1$
DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.087 350.087350.087\,350.087 35 $0.9996$
SAM cran 13 0481304813\,04813 048 $0.9992$ 0.105 450.105450.105\,450.105 45 $0.9999$
limma bioconductor 13 4511345113\,45113 451 $0.9993$ 0.105 240.105240.105\,240.105 24 $0.9998$
ggplot2 cran 14 4411444114\,44114 441 $0.9995$ 0.074 010.074010.074\,010.074 01 $0.9995$
PERMANOVA cran 21 9602196021\,96021 960 $1$ 0.048 610.048610.048\,610.048 61 $0.999$
edgeR bioconductor 12 9231292312\,92312 923 $0.9991$ 0.103 090.103090.103\,090.103 09 $0.9997$
WGCNA cran 13 9151391513\,91513 915 $0.9994$ 0.051 950.051950.051\,950.051 95 $0.9992$
pymol pypi 14 6041460414\,60414 604 $0.9996$ 0.046 610.046610.046\,610.046 61 $0.9989$
tophat pypi $7642$ $0.9987$ 0.073 770.073770.073\,770.073 77 $0.9994$
vegan cran 10 0121001210\,01210 012 $0.9989$ 0.051 720.051720.051\,720.051 72 $0.9991$
lme4 cran 10 5101051010\,51010 510 $0.999$ 0.037 380.037380.037\,380.037 38 $0.9986$
Popular
GSVA bioconductor $2937$ $0.9969$ 0.001 360.001360.001\,360.001 36 $0.1214$
MAST bioconductor $2091$ $0.995$ 0.001 360.001360.001\,360.001 36 $0.1214$
GSA cran $844$ $0.9868$ 0.001 360.001360.001\,360.001 36 $0.2163$
FSA cran $499$ $0.9782$ 0.001 360.001360.001\,360.001 36 $0.1508$
Boruta cran $441$ $0.9754$ 0.001 360.001360.001\,360.001 36 $0.0968$
MixSIAR cran $377$ $0.9714$ 0.001 380.001380.001\,380.001 38 $0.4265$
AntWeb cran $282$ $0.9616$ 0.001 380.001380.001\,380.001 38 $0.4265$
emoji pypi $161$ $0.9374$ 0.001 380.001380.001\,380.001 38 $0.4356$
eureqa pypi $149$ $0.9331$ 0.001 380.001380.001\,380.001 38 $0.4265$
searchlight pypi $137$ $0.9273$ 0.001 360.001360.001\,360.001 36 $0.259$
pathfindR cran $133$ $0.9259$ 0.001 370.001370.001\,370.001 37 $0.4023$
ChIPXpress bioconductor $125$ $0.9213$ 0.001 380.001380.001\,380.001 38 $0.4129$
pypet pypi $125$ $0.9213$ 0.001 360.001360.001\,360.001 36 00
ADAPTS cran $125$ $0.9213$ 0.001 370.001370.001\,370.001 37 $0.3452$
“Nebraska”
calculate_expression pypi $3$ $0.3287$ 0.018 720.018720.018\,720.018 72 $0.9956$
zscore pypi $5$ $0.4326$ 0.008 510.008510.008\,510.008 51 $0.9893$
webpages pypi $2$ $0.2516$ 0.007 070.007070.007\,070.007 07 $0.9872$
setuptools pypi $6$ $0.4689$ 0.005 720.005720.005\,720.005 72 $0.9825$
gpviz pypi $1$ $0.1232$ 0.004 530.004530.004\,530.004 53 $0.9759$
glib pypi $3$ $0.3287$ 0.003 970.003970.003\,970.003 97 $0.9711$
ego pypi $5$ $0.4326$ 0.003 260.003260.003\,260.003 26 $0.9607$
xtick pypi $1$ $0.1232$ 0.003 090.003090.003\,090.003 09 $0.9575$
pauvre pypi $5$ $0.4326$ 0.002 920.002920.002\,920.002 92 $0.9531$
mappy pypi $4$ $0.3863$ 0.002 920.002920.002\,920.002 92 $0.9531$
rle cran $2$ $0.2516$ 0.002 870.002870.002\,870.002 87 $0.9522$
LineagePulse bioconductor $1$ $0.1232$ $0.0027$ $0.9474$
Table 5: Software quadrants, weighted graph (LCC component only)
Software Ecosystem Mentions Centrality
Count Percentile Value Percentile
“Pasteur”
PRISMA cran 15 7971579715\,79715 797 $0.9996$ 0.257 560.257560.257\,560.257 56 $1$
DESeq2 bioconductor 21 2572125721\,25721 257 $0.9999$ 0.174 920.174920.174\,920.174 92 $0.9995$
SAM cran 13 0481304813\,04813 048 $0.999$ 0.211 160.211160.211\,160.211 16 $0.9999$
limma bioconductor 13 4511345113\,45113 451 $0.9991$ 0.210 740.210740.210\,740.210 74 $0.9998$
ggplot2 cran 14 4411444114\,44114 441 $0.9994$ $0.1482$ $0.9994$
PERMANOVA cran 21 9602196021\,96021 960 $1$ 0.097 350.097350.097\,350.097 35 $0.9988$
edgeR bioconductor 12 9231292312\,92312 923 $0.9989$ 0.206 430.206430.206\,430.206 43 $0.9996$
WGCNA cran 13 9151391513\,91513 915 $0.9993$ 0.104 020.104020.104\,020.104 02 $0.999$
pymol pypi 14 6041460414\,60414 604 $0.9995$ 0.093 330.093330.093\,330.093 33 $0.9986$
tophat pypi $7642$ $0.9984$ 0.147 730.147730.147\,730.147 73 $0.9993$
vegan cran 10 0121001210\,01210 012 $0.9986$ 0.103 570.103570.103\,570.103 57 $0.9989$
lme4 cran 10 5101051010\,51010 510 $0.9988$ 0.074 850.074850.074\,850.074 85 $0.9983$
Popular
GSVA bioconductor $2937$ $0.996$ 0.002 720.002720.002\,720.002 72 $0.1185$
MAST bioconductor $2091$ $0.9937$ 0.002 720.002720.002\,720.002 72 $0.1185$
GSA cran $844$ $0.9833$ 0.002 720.002720.002\,720.002 72 $0.2167$
FSA cran $499$ $0.9724$ 0.002 720.002720.002\,720.002 72 $0.148$
Boruta cran $441$ $0.969$ 0.002 720.002720.002\,720.002 72 $0.0923$
MixSIAR cran $377$ $0.9639$ 0.002 760.002760.002\,760.002 76 $0.3878$
AntWeb cran $282$ $0.9516$ 0.002 760.002760.002\,760.002 76 $0.3878$
emoji pypi $161$ $0.9215$ 0.002 760.002760.002\,760.002 76 $0.397$
searchlight pypi $137$ $0.9091$ 0.002 730.002730.002\,730.002 73 $0.254$
pathfindR cran $133$ $0.9073$ 0.002 750.002750.002\,750.002 75 $0.3677$
ChIPXpress bioconductor $125$ $0.9018$ 0.002 760.002760.002\,760.002 76 $0.3765$
pypet pypi $125$ $0.9018$ 0.002 720.002720.002\,720.002 72 00
“Nebraska”
calculate_expression pypi $3$ $0.2982$ $0.0375$ $0.9944$
zscore pypi $5$ $0.3883$ 0.017 050.017050.017\,050.017 05 $0.9865$
setuptools pypi $6$ $0.4215$ 0.011 460.011460.011\,460.011 46 $0.978$
newick pypi $9$ $0.4957$ 0.008 060.008060.008\,060.008 06 $0.9649$
xtick pypi $1$ $0.1305$ 0.006 180.006180.006\,180.006 18 $0.9469$
nanolyse pypi $9$ $0.4957$ 0.005 950.005950.005\,950.005 95 $0.9443$
splatter pypi $8$ $0.4742$ 0.005 840.005840.005\,840.005 84 $0.9422$
pauvre pypi $5$ $0.3883$ 0.005 840.005840.005\,840.005 84 $0.9413$
mappy pypi $4$ $0.3474$ 0.005 840.005840.005\,840.005 84 $0.9413$
rle cran $2$ $0.232$ 0.005 740.005740.005\,740.005 74 $0.9402$
scone pypi $8$ $0.4742$ 0.005 740.005740.005\,740.005 74 $0.9398$
LineagePulse bioconductor $1$ $0.1305$ $0.0054$ $0.9344$
@@ 4 Discussion
The absence of cycles in the network of open-source software packages used in our sample of biomedical science is quite interesting. It suggests a more robust design structure in the universe of scientific software than in the general software world. The lack of cycles also improves our analysis with Katz centrality, as the absence of loops excludes feedback mechanisms that can artificially inflate the centrality of packages on a loop. Packages with low mentions and high Katz centrality are therefore even more critical, as they represent essential dependencies of packages that enable large volumes of research without receiving direct attention. These are the “Nebraskan” packages referenced by [1] that we set out to find. Figure 4 and Tables 3, 4, and 5 provide examples of these packages. As expected, most of these software packages are libraries used by other packages: vctrs [33] (a package for type manipulation in R), withr [34] (a set of tools for safely running functions that change global variables), isoband [35] (a package to generate isolines from gridded data), pauvre [36] (a package for plotting Oxford Nanopore data), newick [37] (read and write Newick data formatted files), setuptools [38] (a packaging library for Python projects). It is interesting that the prominent R “Nebraska” packages belong to the tidyverse system [39], and thus are co-authored by Hadley Wickham.
The difference between the unweighted and weighted versions of our analysis, in Figure 4 can help us parse different subtypes of these outliers. For example, the velvet [40] package is central solely because it is mentioned by one highly cited paper (and dozens of less highly cited ones). This strategy is not central in the unweighted version of the analysis, since citation counts are ignored. Central packages under this representation, therefore, also need to have other packages depend on them and cannot solely rely on mentions from influential papers.
It is interesting that cran and bioconductor dominate the “Nebraska” quadrant for the unweighted graph, while for the weighted graph, pypi is overrepresented. The reason for this difference may warrant additional research.
Altogether, our paper tries to quantify impact as importance and to provide a set of metrics behind the intuition. Of course, this is by itself a perilous process: there could be various flavors of both impact and importance, and people may reasonably disagree about the details. Our contribution is to propose a certain quantification, which agrees with our intuitive understanding.
In addition to possible problems with the exact definition of importance, our work has a number of limitations.
@@
    The network analysis is limited by the quality of disambiguation and linking in the CZ Software Mentions dataset (see the discussion in [11]). Homonyms (different software packages with the same or close names) may change the network statistics, and preliminary research points to considerable potential for incorrect linking in the CZ Software Mentions dataset due to homonyms [7]. Notably, this affects the package velvet (see Figure 3a), which has been falsely linked to a Python package that provides signal processing and communications algorithms in Python [41]. Meanwhile, all of the mentions of “velvet” point to different versions of algorithms for de novo short-read assembly using de Bruijn graphs in genomics, as described in [42] and implemented mostly in C [40]. While this does not change its centrality and our interpretation, the linking error led to its mislabelling as a PyPI package for the visualization in Figure 3.
    Sometimes an open source package also shares a name with a proprietary software (e. g. PRISM, PACE), which leads to confusion and incorrect mention counts.
@@
    Our approach is limited to the open-source packages available in one of the chosen ecosystems. This excludes proprietary software like Excel or GraphPad Prism, and open-source non-package software like Gephi [43] (a complied application). The number of missing packages may be estimated from Table 14 of [11], which estimates the coverage for GitHub at 64.39%, the coverage for CRAN at 8.36%, the coverage for PyPI at 5.86%, and the coverage for Bioconductor at 3.23%. GitHub coverage can be used as a proxy for the share of the open source software: while some packages—perhaps growing in number—are not hosted by GitHub, their fraction remains low [44]. If we assume that each package exists only in one ecosystem (there are packages present in several, but we neglect this effect), the combined coverage by CRAN, PyPI, and Bioconductor is 17.45%, i.e., between a quarter and one third of the open source packages. It is very likely that each of these non-package pieces of software do depend on packages in our focus ecosystems. For open-source non-package software, it may be feasible to identify source repository URLs and then identify packages depended on by the software. That would need to be done by directly inspecting the source code, possibly using Software Bill of Materials (SBOM) tools. On the other hand, some non-package software does not have source code available (including proprietary GUI software such as SPSS, and cloud-based services). Nonetheless, these pieces of code may very well be based on packages in our focal ecosystem (especially if the latter use non-restrictive licenses such as BSD, MIT). In the long term, there is a possibility that the requirement to provide SBOMs for U.S. federal government purchasers111as per Executive Order 14028: Improving the Nation’s Cybersecurity from May 12, 2021 will give insight into the packages on which proprietary software relies [45].
@@
    In our dependency network, we always used the latest version of any package. In reality, dependencies change between the versions. A study of development logs often reveals messages like “deleted the dependence on XYZ” or “added the dependence on ABC”. These changes often make the work of package management software very difficult. A software dependency graph is a living network, constantly evolving, with links added and deleted. In this study, we effectively collapse the time, making an (imperfect) snapshot of the long movie.
@@
    Our method captures the off-the-shelf software packages used for each article. It does not capture ad hoc software written specifically for the given paper, which may be contained in the code accompanying the papers. This code may load software libraries and thus add software dependencies that are absent from our graph.
@@
    The dependencies reflected in the package metadata do not necessarily reflect the actual usage of the software. A library may be loaded, but not used for the actual execution by a given paper. A more reliable method would be to reproduce the analysis done in each paper and capture the actual library calls—which is probably prohibitively difficult for the number of papers covered.
@@
    Many Python and R packages ultimately depend on libraries outside their ecosystems (most often C and Fortran libraries). These dependencies are not easily captured by the package management infrastructure and thus are outside of our analysis.
@@
    Our dependency network does not include the concept of alternatives, when package $A$ depends on the functionality that can be provided either by package $B_{1}$, or package $B_{2}$. Some package managers like Debian’s apt have the possibility of specifying alternatives, but this is not a common feature of package managers.
@@
    Different communities place different emphasis on infrastructure elements such as tests. This may lead to the over-representation of testing infrastructure packages in some cases, for example, in the Python ecosystem. We do not try to attempt to decide which elements of the infrastructure are “actually important”; instead, we rely on the judgment of the community of users and developers while recognizing that such judgment may be different across the landscape.
@@
    There are different software package repositories and ecosystems for Python and R with greatly overlapping dependency graphs. We used PyPI, CRAN and Bioconductor for this study. Alternative repositories that provide packages for Python and R, such as Conda channels (conda-forge, bioconda, etc.), are also worth investigating.